A 3.215 kg unknown substance was transformed from solid to liquid by condensing 10.57 kg of steam under atmospheric condition


What is the initial temperature of steam in degF?

How much heat was gained by the steam in kJ?

What is the heat of fusion of the unknown substance in kj/g?


1
Expert's answer
2021-11-30T21:29:01-0500

1) 10.57/3.215=3.3•

2) 10.57×3.3=34.75

3) 3.215×3.3=10.6095